# Deep Learning of Multivariate Extremes via a Geometric Representation

### v1

@@ Abstract
The study of geometric extremes, where extremal dependence properties are inferred from the deterministic limiting shapes of scaled sample clouds, provides an exciting approach to modelling the extremes of multivariate data. These shapes, termed limit sets, link together several popular extremal dependence modelling frameworks. Although the geometric approach is becoming an increasingly popular modelling tool, current inference techniques are limited to a low dimensional setting ($d\leq 3$), and generally require rigid modelling assumptions. In this work, we propose a range of novel theoretical results to aid with the implementation of the geometric extremes framework and introduce the first approach to modelling limit sets using deep learning. By leveraging neural networks, we construct asymptotically-justified yet flexible semi-parametric models for extremal dependence of high-dimensional data. We showcase the efficacy of our deep approach by modelling the complex extremal dependencies between meteorological and oceanographic variables in the North Sea off the coast of the UK.
Keywords: extremal dependence; geometric extremes; limit sets; neural networks
@@ 1 Introduction
Multivariate extreme value theory is a branch of statistics that deals with understanding relationships between the extremes of multiple variables. A wide variety of modelling approaches for multivariate extremes (or, equivalently, extremal dependence) have been proposed; classical approaches use the framework of regular variation (see, e.g., Tawn,, 1990; Rootzén and Tajvidi,, 2006; Einmahl and Segers,, 2009), but these models are restrictive in the forms of extremal dependence that they can capture (Huser et al.,, 2024). In particular, they can only capture asymptotic dependence (Coles et al.,, 1999), where extremes of a random vector occur when all of its components are jointly large. Assuming this form of extremal dependence for data is unrealistic in many applications, and numerous works have advocated against the use of regular variation models for environmental studies (Opitz,, 2016; Dawkins and Stephenson,, 2018; Huser et al.,, 2024).
The first approach to modelling non-asymptotically dependent data was provided by Ledford and Tawn, (1996, 1997); see, also, hidden regular variation (Resnick,, 2002). Letting $\bm{X}_{E}:=(X_{E,1},\dots,X_{E,d})$ denote a random vector with standard exponential margins, Ledford and Tawn, (1996) assume that, as $u\to\infty$,
@@
as $u\to\infty$, where $L(\cdot\;;\bm{w})$ is a slowly varying function and $\lambda(\cdot)$ denotes the ADF; the latter provides information about the joint tail of $\bm{X}_{E}$, and we have $\eta=\{\sqrt{d}\lambda(d^{-1/2},\ldots,d^{-1/2})\}^{-1}$. Model (2) can capture both extremal dependence regimes, with asymptotic dependence implying the lower bound, $\lambda(\bm{w})=\max(\bm{w})$ for all $\bm{w}\in\mathcal{S}^{d-1}_{+}$. Loosely speaking, the angle $\bm{w}$ is the direction in $\mathbb{R}_{+}^{d}$ for which the joint tail region in (2) is defined. This model has been successfully applied in environmental applications by, e.g., Murphy-Barltrop et al., (2023), Murphy-Barltrop and Wadsworth, (2024), and Murphy-Barltrop et al., 2024b .
@@
Several of the aforementioned models introduced are defined for random vectors exhibiting standard margins with finite lower bounds, e.g., Pareto, Fréchet, or exponential, which limits the study of extremal dependence. In particular, when applied to random vectors with double tailed margins, such as Laplace, these modelling approaches reduce the study of extremal dependence to data observed only in the positive orthant, +$\mathbb{R}^{d}_{+}$, of $\mathbb{R}^{d}$; see Section 2 for further discussion. This can be restrictive in practical applications where different regions of low joint probability mass may be of interest. Consequently, many recent works have introduced modelling approaches for data on standard Laplace margins. For example, Keef et al., (2013) extended the model of Heffernan and Tawn, (2004) to Laplace margins, with the resulting framework providing greater flexibility and interpretability. Moreover, Mackay and Jonathan, (2023), Papastathopoulos et al., (2024) and Murphy-Barltrop et al., 2024a demonstrate that models on Laplace margins permit evaluation of the joint tail behaviour of random vectors in all $2^{d}$ orthants of $\mathbb{R}^{d}$. To demonstrate this, Figure 1 illustrates extreme density contours for a bivariate Gaussian copula with correlation parameter $\rho=-0.5$ on both standard exponential and standard Laplace margins. Hereafter, we use $\bm{X}:=(X_{1},\dots,X_{d})^{T}$ to denote a $d$-dimensional random vector with standard Laplace margins, with distribution function $F_{\bm{X}}(\cdot)$ and continuous density function $f_{\bm{X}}(\cdot)$.
@@
Figure 1: Density contours for a bivariate random vector $\bm{X}$ with a Gaussian copula (with negative correlation) and with standard exponential (left) or standard Laplace (right) margins. The density levels for each colour are given in the legend in the left panel.
@@
Recent theoretical developments for multivariate extremes have focused on geometric extremes, whereby extremal dependence properties of $\bm{X}$ can be inferred directly from the deterministic limiting shapes of scaled sample clouds. Let $C_{n}:={\{\bm{X}_{i}/r_{n}\}^{n}_{i=1}}$ denote $n$ independent copies of $\bm{X}$, scaled by a suitably chosen positive sequence $(r_{n})_{n\in\mathbb{N}}$ satisfying $r_{n}\to\infty$ as $n\to\infty$. Under mild conditions, $C_{n}$ converges in probability, with respect to the Hausdorff metric, onto the compact, star-shaped limit set $\mathcal{G}:=\{\bm{x}:g(\bm{x})\leq 1\}\subset[-1,1]^{d}$, where $g:\mathbb{R}^{d}\mapsto\mathbb{R}_{+}$ is the gauge function of $\mathcal{G}$ (Fisher,, 1969; Davis et al.,, 1988; Kinoshita and Resnick,, 1991). A sufficient condition for convergence onto $\mathcal{G}$ is that
@@
or, equivalently, $g(\bm{x})=\lim_{t\to\infty}[-\log f_{\bm{X}}(t\bm{x})]/t$ (Balkema and Nolde,, 2010). Following Nolde, (2014), we also define the unit-level, or boundary, set $\partial\mathcal{G}:=\{\bm{x}:g(\bm{x})=1\}\subset\mathcal{G}$. For standard Laplace margins, it suffices to set $r_{n}=\log(n/2)$ to achieve the required convergence (Papastathopoulos et al.,, 2024). To demonstrate this, Figure 2 illustrates the limit set $\mathcal{G}$ and unit-level set $\partial\mathcal{G}$ for three copulas, alongside $n=10,000$ samples $\{\bm{x}_{i}/\log(n/2)\}_{i=1}^{n}$ from each copula; formal definitions of these copulas are given in Section 4. One can observe that the (finite) observed sample clouds lie approximately within the theoretical limit set. Hereafter, we implicitly assume that any $\bm{X}$ satisfies the conditions for convergence onto $\mathcal{G}$.
@@
Figure 2: Scaled sample clouds of size $n=10,000$ from Gaussian (left), Student-t (middle), and logistic (right) copulas on standard Laplace margins. For each panel, the shaded regions and solid red lines give the limit set $\mathcal{G}$ and the unit-level set $\partial\mathcal{G}$, respectively, while the solid blue lines denote the set $\{\bm{w}/\Lambda(\bm{w}):\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}\}$; see Section 2.4 for further details. The black dotted lines denote the unit square.
@@
Nolde, (2014) and Nolde and Wadsworth, (2022) illustrate that the shape of $\partial\mathcal{G}$, or equivalently $\mathcal{G}$, is directly related to extremal dependence of $\bm{X}$. Specifically, $\partial\mathcal{G}$ links together several representations for multivariate extremes: from $\partial\mathcal{G}$, we can determine all parameters associated with the models described in equations (1) and (2), as well as those proposed by Heffernan and Tawn, (2004) and Simpson et al., (2020). Taking, e.g., the ADF, we have that $\lambda(\bm{w})=||\bm{w}||_{\infty}\times\mathfrak{r}_{\bm{w}}^{-1}$ for all $\bm{w}\in\mathcal{S}^{d-1}_{+}$, where $\mathfrak{r}_{\bm{w}}:=\max\{\mathfrak{r}\in[0,1]:\mathfrak{r}\mathcal{R}_{\bm{w}}\cap\partial\mathcal{G}\neq\emptyset\}$ is a $\bm{w}$-dependent coefficient used to scale back the set $\mathcal{R}_{\bm{w}}:=\bigotimes_{i=1,\ldots,d}[w_{i}/||\bm{w}||_{\infty},\infty]$ to intersect with $\partial\mathcal{G}$. Such parameters allow us to quantify extremal dependence of $\bm{X}$ in an interpretable manner. Furthermore, from a practical perspective, estimates of $\partial\mathcal{G}$ can be used to estimate extreme statistics; for example, joint tail probabilities (Wadsworth and Campbell,, 2024), return curves (Murphy-Barltrop et al., 2024b, ), and return level sets (Papastathopoulos et al.,, 2024). The unit-level set $\partial\mathcal{G}$ therefore offers a high degree of practical utility for inference of multivariate extremes, and thus its accurate estimation is of particular importance.
@@
Owing to this perspective, recent works have introduced techniques for the estimation of $\partial\mathcal{G}$, on both standard exponential and standard Laplace margins. For the former margin, Simpson and Tawn, 2024a and Majumder et al., 2024b proposed semi-parametric techniques using generalised additive models (GAMs) and Bézier curves, respectively, to approximate $\partial\mathcal{G}$, while Wadsworth and Campbell, (2024) used a parametric-copula-based model. For Laplace margins, Papastathopoulos et al., (2024) proposed a latent Gaussian model to approximate the shape of $\partial\mathcal{G}$, while Murphy-Barltrop et al., 2024a used GAMs to estimate $\partial\mathcal{G}$ using the model introduced by Mackay and Jonathan, (2023). Within these approaches, a wide range of statistical techniques have been considered in both Bayesian and frequentist settings, and it is possible to obtain accurate estimates of $\partial\mathcal{G}$ for a wide range of dependence structures. However, current techniques for modelling and estimating $\partial\mathcal{G}$ have several shortcomings: i) all existing approaches are limited to a low-dimensional setting ($d\leq 3$), ii) the restriction of several approaches to standard exponential margins offers a limited perspective for evaluating joint tail properties (see Figure 1), and iii) one must always specify parametric or semi-parametric forms for quantities related to $\partial\mathcal{G}$, or select a large number of tuning parameters. Therefore, the existing estimation techniques for $\partial\mathcal{G}$ offer limited practical utility, motivating novel developments.
@@
Recent literature combining extreme value theory and deep learning has seen the construction of flexible, computationally-scalable models for univariate extremes (see, e.g., Pasche and Engelke,, 2022; Richards and Huser,, 2022; Cisneros et al.,, 2024) and generative models for multivariate and spatial extremes (Boulaguiem et al.,, 2022; Lafon et al.,, 2023; Zhang et al.,, 2023; Majumder et al., 2024a, ). While Hasan et al., (2022) used neural networks to build flexible models for asymptotically-dependent multivariate extremes, deep learning is yet to be exploited in the construction of models that can capture non-asymptotically dependent data structures. Here, we propose the first deep learning-based approach for modelling $\partial\mathcal{G}$, referred to hereafter as the DeepGauge framework. Our approach uses neural networks to perform inference on multivariate extremes, and represents a significant step towards flexible and robust models that require minimal parametric assumptions. The use of deep learning methods gives the DeepGauge framework a high degree of flexibility and, as we demonstrate in Sections 4 and 5, allows us to capture a wide variety of extremal dependence structures. Furthermore, it can be applied in higher dimensional settings ($d\geq 4$) than existing techniques for estimating $\partial\mathcal{G}$, and also requires fewer modelling assumptions.
@@
The paper is organised as follows. Section 2 outlines the theory underpinning the DeepGauge framework, alongside novel theoretical results pertaining to the unit-level set $\partial\mathcal{G}$ and its estimation. Section 3 outlines our methodology for estimating $\partial\mathcal{G}$ and related quantities, with Section 3.3 detailing our neural network-based representation for geometric extremes and Section 3.5 introducing novel diagnostic tools for validating model fits in high dimensional settings. Section 4 provides a simulation study showcasing the efficacy of our framework for inferring the extremal dependence of random vectors. Section 5 provides an application to the NORA10 hindcast data set of meteorological and oceanographic (metocean) variables in the North Sea that exhibit complex dependence structures. We conclude in Section 6 with a discussion and outlook on future work.
@@ 2.1 Overview of the angular-radial decomposition
To estimate the unit-level set $\partial\mathcal{G}$ and its corresponding gauge function $g(\cdot)$, we first decompose $\bm{X}$ into angular and radial components, and then model the radii conditional on a fixed angle. While one could select one of many radial-angular systems for this decomposition, we follow advocacy by Murphy-Barltrop et al., 2024a and define angular and radial components via the Euclidean norm. For any $\bm{X}\in\mathbb{R}^{d}\setminus\bm{0}_{d}$, with $\bm{0}_{d}:=(0,\ldots,0)^{T}$, define $(R,\bm{W})$ by $\bm{X}\mapsto(R,\bm{W}):=(\|\bm{X}\|,\bm{X}/\|\bm{X}\|)$ for $R>0$ and $\bm{W}\in\mathcal{S}^{d-1}$, where $\mathcal{S}^{d-1}:=\{\bm{x}\in\mathbb{R}^{d}:||\bm{x}||=1\}$ denotes the unit $(d-1)$-sphere. It follows that $\bm{X}=R\bm{W}$, which implies that $\bm{X}$ is completely determined by the joint behaviour of $R$ and $\bm{W}$. It is trivial to show that mapping $t:\mathbb{R}^{d}\setminus\bm{0}_{d}\mapsto\mathbb{R}_{+}\times\mathcal{S}^{d-1}$, where $t(\bm{x}):=(||\bm{x}||,\bm{x}/||\bm{x}||)$, is bijective; thus, no information is lost through considering $(R,\bm{W})$. Loosely speaking, $R$ is the magnitude of an event, while $\bm{W}$ defines its direction, i.e., in which orthant of $\mathbb{R}^{d}$ the event occurs. Note that for any $d\in\mathbb{N}$ there exist $2^{d}$ possible orthants, as each component can be either positive or negative.
@@
We now recall some theoretical properties of gauge functions and limit sets. The star-shapedness of the limit set $\mathcal{G}$ implies that, for any $\bm{x}\in\mathcal{G}$ and $t>0$, we have $t\bm{x}\in\mathcal{G}$. Moreover, if $\bm{0}_{d}\in\mathcal{G}$ and $g(\bm{0}_{d})<1$, we have that the line segment $\{\bm{0}_{d}+t\bm{x}:t\in[0,1]\}\subset\mathcal{G}$ for any $\bm{x}\in\mathcal{G}$. Furthermore, one can show that the componentwise maxima and minima of $\mathcal{G}$ equal $\bm{1}_{d}$ and $-\bm{1}_{d}$ respectively, implying that $\mathcal{G}$ (and, thus, the unit-level set $\partial\mathcal{G}$) must touch all boundaries of the unit hypercube $[-1,1]^{d}$ at least once. We further note that the gauge function $g(\cdot)$ of $\mathcal{G}$ is 1-homogeneous, i.e., $g(t\bm{x})=tg(\bm{x})$ for any $\bm{x}\in\mathbb{R}^{d},t\in\mathbb{R}_{+}$.
@@
Using the radial-angular decomposition of $\bm{X}$ and the star-shaped property of $\mathcal{G}$, we can reformulate the unit-level set as $\partial\mathcal{G}=\left\{r\bm{w}:r>0,\bm{w}\in\mathcal{S}^{d-1},g(r\bm{w})=1\right\}.$ Homogeneity of $g(\cdot)$ implies that, for any $\bm{w}\in\mathcal{S}^{d-1}$, the radial value of the corresponding point on the unit-level set must be $1/g(\bm{w})$; hence, $\partial\mathcal{G}=\left\{\bm{w}/g(\bm{w}):\bm{w}\in\mathcal{S}^{d-1}\right\}.$ This reformulation has the powerful implication that, to determine $\partial\mathcal{G}$, we need only to evaluate $g(\cdot)$ on $\mathcal{S}^{d-1}$. Illustrations of the radial-angular representations for the sets $\mathcal{G}$ and $\partial\mathcal{G}$ are given in Figure 3.
@@
Figure 3: Gaussian limit set $\mathcal{G}$ (shaded regions) and unit-level set $\partial\mathcal{G}$ (solid red lines) in Cartesian (left) and polar (right) coordinates. The black dotted lines in both plots denote the unit square. For the right panel, we plot the radii of the unit-level set (i.e., $1/g(\bm{w})$) against the standard polar angle, $\cos^{-1}(w_{1})=\sin^{-1}(w_{2})\in[0,2\pi)$ for each $(w_{1},w_{2})\in\mathcal{S}^{1}$.

### v2

@@ Abstract
The study of geometric extremes, where extremal dependence properties are inferred from the deterministic limiting shapes of scaled sample clouds, provides an exciting approach to modelling the extremes of multivariate data. These shapes, termed limit sets, link several popular extremal dependence modelling frameworks. Although the geometric approach is becoming an increasingly popular modelling tool for multivariate extremes, current inference techniques are limited to a low dimensional setting ($d\leq 4$), and generally require rigid modelling assumptions. In this work, we propose a range of novel theoretical results to aid with the implementation of the geometric extremes framework and introduce the first approach to modelling limit sets using deep learning. By leveraging neural networks, we construct asymptotically-justified and flexible semi-parametric models for extremal dependence of high-dimensional data. We showcase the efficacy of our approach by modelling the complex extremal dependencies between meteorological and oceanographic variables in the North Sea.
Keywords: extremal dependence; geometric extremes; neural networks; deep learning
@@ 1 Introduction
Multivariate extreme value theory seeks to understand relationships between the extremes of multiple variables. A wide variety of modelling approaches for multivariate extremes (or, equivalently, extremal dependence) have been proposed; classical approaches use the framework of regular variation (see, e.g., Tawn,, 1990; Rootzén and Tajvidi,, 2006; Einmahl and Segers,, 2009), but these models are restrictive in the forms of extremal dependence that they can capture (Huser et al.,, 2024). In particular, they can only provide accurate inference for random vectors exhibiting asymptotic dependence, i.e., when all variables are simultaneously extreme (Ledford and Tawn,, 1997; Heffernan and Tawn,, 2004). Assuming this form of extremal dependence for data is unrealistic in many applications, and numerous works have advocated against the use of regular variation models for environmental studies (e.g., Opitz,, 2016; Dawkins and Stephenson,, 2018; Huser et al.,, 2024).
The first approach to modelling asymptotically independent data was provided by Ledford and Tawn, (1996, 1997); see, also, hidden regular variation (Resnick,, 2002). Letting $\bm{X}_{E}:=(X_{E,1},\dots,X_{E,d})^{T}$ denote a random vector with standard exponential margins, Ledford and Tawn, (1996) assumed that, as $u\to\infty$,
@@
as $u\to\infty$, where $L(\cdot\;;\bm{w})$ is a slowly varying function and $\lambda(\cdot)$ denotes the ADF; the latter provides information about the joint tail of $\bm{X}_{E}$, and we have $\eta=\{\sqrt{d}\lambda(d^{-1/2},\ldots,d^{-1/2})\}^{-1}$. Model (2) can capture both extremal dependence regimes, with asymptotic dependence implying the lower bound, $\lambda(\bm{w})=\max(\bm{w})$, for all $\bm{w}\in\mathcal{S}^{d-1}_{+}$. Loosely speaking, the angle $\bm{w}$ is the direction in $\mathbb{R}_{+}^{d}$ for which the joint tail region in (2) is defined. This model has been successfully applied in environmental applications by, e.g., Murphy-Barltrop et al., (2023), Murphy-Barltrop and Wadsworth, (2024), and Murphy-Barltrop et al., 2024b .
@@
Several of the aforementioned models introduced are defined for random vectors exhibiting standard margins with finite lower bounds, e.g., Pareto, Fréchet, or exponential, which limits the study of extremal dependence. In particular, taking random vectors with double tailed margins and applying such approaches (following marginal transformation), these models reduce the study of extremal dependence to data observed only in the positive orthant, +$\mathbb{R}^{d}_{+}$, of $\mathbb{R}^{d}$; see Section 2 for further discussion. This can be restrictive in practical applications where different regions of low joint probability mass may be of interest. Consequently, many recent works have introduced modelling approaches for data on standard Laplace margins. For example, Keef et al., (2013) extended the model of Heffernan and Tawn, (2004) to Laplace margins, with the resulting framework providing greater flexibility and interpretability. Moreover, Mackay and Jonathan, (2023), Simpson and Tawn, 2024b , Papastathopoulos et al., (2024) and Murphy-Barltrop et al., 2024a demonstrate that models on Laplace margins permit evaluation of the joint tail behaviour of random vectors in all $2^{d}$ orthants of $\mathbb{R}^{d}$. To demonstrate this, Figure 1 illustrates extreme isodensity contours for a bivariate Gaussian copula with correlation parameter $\rho=-0.5$ on both standard exponential and standard Laplace margins. One can observe that the Laplace representations offers a more detailed perspective of the entire dependence structure. Hereafter, we use $\bm{X}:=(X_{1},\dots,X_{d})^{T}$ to denote a $d$-dimensional random vector with standard Laplace margins, with distribution function $F_{\bm{X}}(\cdot)$ and continuous density function $f_{\bm{X}}(\cdot)$.
@@
Figure 1: Isodensity contours for a bivariate random vector $\bm{X}$ with a Gaussian copula (with negative correlation) and with standard exponential (left) or standard Laplace (right) margins. The density levels for each colour are given in the legend in the left panel.
@@
Recent theoretical developments for multivariate extremes have focused on geometric extremes, whereby extremal dependence properties of $\bm{X}$ can be inferred directly from the deterministic limiting shapes of scaled sample clouds. Let $C_{n}:={\{\bm{X}_{i}/r_{n}\}^{n}_{i=1}}$ denote $n$ independent copies of $\bm{X}$, scaled by a suitably chosen positive sequence $(r_{n})_{n\in\mathbb{N}}$ satisfying $r_{n}\to\infty$ as $n\to\infty$. Under mild conditions, $C_{n}$ converges in probability, with respect to the Hausdorff metric, onto the compact, star-shaped limit set $\mathcal{G}:=\{\bm{x}:g(\bm{x})\leq 1\}\subset[-1,1]^{d}$, where $g:\mathbb{R}^{d}\mapsto\mathbb{R}_{+}$ is the gauge function of $\mathcal{G}$ (Fisher,, 1969; Davis et al.,, 1988; Kinoshita and Resnick,, 1991). We formally define star-shaped sets in Section 2.1 below. A sufficient condition for convergence onto $\mathcal{G}$ is that
@@
or, equivalently, $g(\bm{x})=\lim_{t\to\infty}[-\log f_{\bm{X}}(t\bm{x})]/t$ (Balkema and Nolde,, 2010). Following Nolde, (2014), we also define the unit-level, or boundary, set $\partial\mathcal{G}:=\{\bm{x}:g(\bm{x})=1\}\subset\mathcal{G}$. For standard Laplace margins, it suffices to set $r_{n}=\log(n/2)$ to achieve the required convergence (Papastathopoulos et al.,, 2024). To demonstrate this convergence, Figure 2 illustrates the limit set $\mathcal{G}$ and unit-level set $\partial\mathcal{G}$ for three copulas, alongside $n=10,000$ samples $\{\bm{x}_{i}/\log(n/2)\}_{i=1}^{n}$ from each copula; formal definitions of these copulas are given in Section 4. One can observe that the (finite) observed sample clouds lie approximately within the theoretical limit set. Hereafter, we implicitly assume that any $\bm{X}$ satisfies the conditions for convergence onto $\mathcal{G}$.
@@
Figure 2: Scaled sample clouds of size $n=10,000$ from Gaussian (left), Student-t (middle), and logistic (right) copulas on standard Laplace margins. For each panel, the shaded regions and solid red lines give the limit set $\mathcal{G}$ and the unit-level set $\partial\mathcal{G}$, respectively, while the solid blue lines denote the set $\{\bm{w}/\Lambda(\bm{w}):\bm{w}\in\mathcal{S}^{d-1}\setminus\mathcal{A}\}$; see Section 2.4 for further details. The black dashed lines denote the unit square.
@@
Nolde, (2014) and Nolde and Wadsworth, (2022) show that the shape of $\partial\mathcal{G}$, or equivalently $\mathcal{G}$, is directly related to extremal dependence of $\bm{X}$. Specifically, $\partial\mathcal{G}$ links several representations for multivariate extremes: from $\partial\mathcal{G}$, we can immediately determine parameters of the models described in equations (1) and (2), as well as those proposed by Heffernan and Tawn, (2004) and Simpson et al., (2020). Taking, e.g., the ADF, we have that $\lambda(\bm{w})=||\bm{w}||_{\infty}\times\mathfrak{r}_{\bm{w}}^{-1}$ for all $\bm{w}\in\mathcal{S}^{d-1}_{+}$, where $\mathfrak{r}_{\bm{w}}:=\max\{\mathfrak{r}\in[0,1]:\mathfrak{r}\mathcal{R}_{\bm{w}}\cap\partial\mathcal{G}\neq\emptyset\}$ is a $\bm{w}$-dependent coefficient used to scale the set $\mathcal{R}_{\bm{w}}:=\bigotimes_{i=1,\ldots,d}[w_{i}/||\bm{w}||_{\infty},\infty]$ to intersect with $\partial\mathcal{G}$. Such parameters provide an interpretable quantification of extremal dependence. Furthermore, from a practical perspective, estimates of $\partial\mathcal{G}$ can be used to estimate extreme statistics, for example, joint tail probabilities (Wadsworth and Campbell,, 2024), return curves (Murphy-Barltrop et al., 2024b, ), and return level sets (Papastathopoulos et al.,, 2024). The unit-level set $\partial\mathcal{G}$ therefore offers a high degree of practical utility for inference of multivariate extremes, and thus its accurate estimation is of particular importance.
@@
Owing to this perspective, recent works have introduced techniques for the estimation of $\partial\mathcal{G}$, on both standard exponential and standard Laplace margins. For the former margin, Simpson and Tawn, 2024a and Majumder et al., 2024b proposed semi-parametric techniques using generalised additive models (GAMs) and Bézier splines, respectively, to approximate $\partial\mathcal{G}$, while Wadsworth and Campbell, (2024) used a parametric copula-based model. For Laplace margins, Papastathopoulos et al., (2024) proposed a latent Gaussian model to approximate the shape of $\partial\mathcal{G}$, while Murphy-Barltrop et al., 2024a used GAMs to estimate $\partial\mathcal{G}$ using the model introduced by Mackay and Jonathan, (2023). Within these approaches, a wide range of statistical techniques have been considered in both Bayesian and frequentist settings, and it is possible to obtain accurate estimates of $\partial\mathcal{G}$ for a large variety of dependence structures. However, current techniques for modelling and estimating $\partial\mathcal{G}$ have several shortcomings: i) all existing approaches are limited to a low-dimensional setting ($d\leq 4$), ii) the restriction of several approaches to standard exponential margins offers a limited perspective for evaluating joint tail properties (see Figure 1), iii) one must always specify parametric or semi-parametric forms for quantities related to $\partial\mathcal{G}$, or select a large number of tuning parameters, and iv) the existing techniques for Laplace margins do not guarantee the theoretical properties of limit sets. Therefore, the existing estimation techniques for $\partial\mathcal{G}$ offer limited practical utility, motivating novel developments.
@@
Recent literature combining extreme value theory and deep learning has seen the construction of flexible, computationally-scalable models for univariate extremes (see, e.g., Pasche and Engelke,, 2022; Richards and Huser,, 2022; Cisneros et al.,, 2024), generative models for multivariate and spatial extremes (Boulaguiem et al.,, 2022; Lafon et al.,, 2023; Zhang et al.,, 2023; Majumder et al., 2024a, ), and classifiers for extremal dependence (Ahmed et al.,, 2022; Wixson and Cooley,, 2024). While Hasan et al., (2022) used neural networks to build flexible models for asymptotically-dependent multivariate extremes, deep learning is yet to be exploited in the construction of models that can capture non-asymptotically dependent data structures. Here, we propose the first deep learning-based approach for modelling $\partial\mathcal{G}$, referred to hereafter as the DeepGauge framework. The use of deep learning methods gives the DeepGauge framework a high degree of flexibility and, as we demonstrate in Sections 4 and 5, allows us to capture a wide variety of extremal dependence structures. Furthermore, it can be applied in higher dimensional settings ($d>4$) than existing techniques for estimating $\partial\mathcal{G}$, and also requires fewer modelling assumptions.
@@
The paper is organised as follows. Section 2 outlines the theory underpinning the DeepGauge framework, alongside novel theoretical results pertaining to the unit-level set $\partial\mathcal{G}$ and its estimation for standard Laplace margins. Section 3 outlines our methodology for estimating $\partial\mathcal{G}$ and related quantities, with Section 3.3 detailing our neural network-based representation for geometric extremes and Section 3.5 introducing diagnostics for validating model fits in high dimensional settings. Section 4 provides a simulation study showcasing the efficacy of our framework for inferring the extremal dependence of random vectors. Section 5 provides an application to the NORA10 hindcast dataset of meteorological and oceanographic (metocean) variables in the North Sea that exhibit complex dependence structures. We conclude in Section 6 with a discussion and outlook on future work.
@@ 2.1 Overview of the angular-radial decomposition
To estimate the unit-level set $\partial\mathcal{G}$ and its corresponding gauge function $g(\cdot)$, we first decompose $\bm{X}$ into angular and radial components, and then model the radii conditional on a fixed angle. While one could select one of many radial-angular systems for this decomposition, we follow Murphy-Barltrop et al., 2024a and define angular and radial components via the Euclidean norm. For any $\bm{X}\in\mathbb{R}^{d}\setminus\bm{0}_{d}$, with $\bm{0}_{d}:=(0,\ldots,0)^{T}$, define $(R,\bm{W})$ by $\bm{X}\mapsto(R,\bm{W}):=(\|\bm{X}\|,\bm{X}/\|\bm{X}\|)$ for $R>0$ and $\bm{W}\in\mathcal{S}^{d-1}$, where $\mathcal{S}^{d-1}:=\{\bm{x}\in\mathbb{R}^{d}:||\bm{x}||=1\}$ denotes the unit $(d-1)$-sphere. It follows that $\bm{X}=R\bm{W}$, implying that $\bm{X}$ is completely determined by the behaviour of $(R,\bm{W})$. It is trivial to show that the mapping $t:\mathbb{R}^{d}\setminus\bm{0}_{d}\mapsto\mathbb{R}_{+}\times\mathcal{S}^{d-1}$, where $t(\bm{x}):=(||\bm{x}||,\bm{x}/||\bm{x}||)$, is bijective; thus, no information is lost through considering $(R,\bm{W})$. Loosely speaking, $R$ is the magnitude of an event, while $\bm{W}$ defines its direction, e.g., in which orthant of $\mathbb{R}^{d}$ the event occurs.
@@
We now recall some theoretical properties of gauge functions and limit sets. The star-shapedness of the limit set $\mathcal{G}$ implies that, for any $\bm{x}\in\mathcal{G}$ and $t>0$, we have $t\bm{x}\in\mathcal{G}$. Moreover, if $\bm{0}_{d}\in\mathcal{G}$ and $g(\bm{0}_{d})<1$, we have that the line segment $\{\bm{0}_{d}+t\bm{x}:t\in[0,1]\}\subset\mathcal{G}$ for any $\bm{x}\in\mathcal{G}$. Furthermore, one can show that the componentwise maxima and minima of $\mathcal{G}$ equal $\bm{1}_{d}$ and $-\bm{1}_{d}$ respectively, implying that $\mathcal{G}$ (and, thus, the unit-level set $\partial\mathcal{G}$) must touch all boundaries of the unit hypercube $[-1,1]^{d}$ at least once. Finally, we note that the gauge function $g(\cdot)$ of $\mathcal{G}$ is 1-homogeneous, i.e., $g(t\bm{x})=tg(\bm{x})$ for any $\bm{x}\in\mathbb{R}^{d},t\in\mathbb{R}_{+}$.
@@
Using the radial-angular decomposition of $\bm{X}$ and the star-shaped property of $\mathcal{G}$, we can reformulate the unit-level set as $\partial\mathcal{G}=\left\{r\bm{w}:r>0,\bm{w}\in\mathcal{S}^{d-1},g(r\bm{w})=1\right\}.$ Homogeneity of $g(\cdot)$ implies that, for any $\bm{w}\in\mathcal{S}^{d-1}$, the radial value of the corresponding point on the unit-level set must be $1/g(\bm{w})$; hence, $\partial\mathcal{G}=\left\{\bm{w}/g(\bm{w}):\bm{w}\in\mathcal{S}^{d-1}\right\}.$ This reformulation has the powerful implication that, to determine $\partial\mathcal{G}$, we need only to evaluate $g(\cdot)$ on $\mathcal{S}^{d-1}$. Illustrations of the radial-angular representations for $\mathcal{G}$ and $\partial\mathcal{G}$ are given in Figure 3.
@@
Figure 3: Gaussian limit set $\mathcal{G}$ (shaded regions) and unit-level set $\partial\mathcal{G}$ (solid red lines) in Cartesian (left) and polar (right) coordinates. The black dashed lines in both plots denote the unit square. For the right panel, we plot the radii of the unit-level set (i.e., $1/g(\bm{w})$) against the standard polar angle, $\cos^{-1}(w_{1})=\sin^{-1}(w_{2})\in[0,2\pi)$ for each $(w_{1},w_{2})\in\mathcal{S}^{1}$.